OpenStudy (melstutes):

angle at peak of roof The house shown below has a symmetrical roof. What is the measure of the angle the roof makes at the peak (angle x)?

OpenStudy (melstutes):

PLEASE Help - I think I would use sin - opposite /hypotenuse. I believe opposite side is 12 and hypotenuse is 15,

OpenStudy (adrimit):

\[\;\sin\;\frac{x}{2}=\frac{12}{15}=\frac{4}{5} \\x=2\arcsin\left(\frac{4}{5}\right)\]